### v2

# Cross-Language Dependencies: An Empirical Study of Kotlin-Java

Abstract.
Background: Since Google introduced Kotlin as an official programming language for developing Android apps in 2017, Kotlin has gained widespread adoption in Android development. The interoperability of Java and Kotlin’s design nature allows them to coexist and interact with each other smoothly within a project. Aims: However, there is limited research on how Java and Kotlin interact with each other in real-world projects and what challenges are faced during these interactions. The answers to these questions are key to understanding these kinds of cross-language software systems. Methods: In this paper, we implemented a tool named DependExtractor, which can extract 11 kinds of Kotlin-Java dependencies, and conducted an empirical study of 23 Kotlin-Java real-world projects with 3,227 Java and 8,630 Kotlin source files. Results: Our findings revealed that Java and Kotlin frequently interact with each other in these cross-language projects, with access and call dependency types being the most dominant. Compared to files interacting with other files in the same language, Java/Kotlin source files, which participate in the cross-language interactions, undergo more commits. Additionally, among all Kotlin-Java problematic interactions, we identified seven common mistakes, along with their fixing strategies. Conclusions: The findings of this study can help developers understand and address the challenges in Kotlin-Java projects.
Kotlin, Java, Code Dependency, Cross Language
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ACM / IEEE International Symposium on Empirical Software Engineering and Measurement (ESEM); October 20–25, 2024; Barcelona, Spainisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Since Google introduced Kotlin as an official programming language for developing Android apps in 2017, Kotlin has gained widespread adoption in Android development. According to recent studies, a significant of Android apps have been continuously migrated from Java to Kotlin (Ardito et al., 2020; Mateus and Martinez, 2020). The interoperability of Java and Kotlin’s design nature allows them to coexist and evolve smoothly in a single project. Meanwhile, understanding the dependencies among software entities is fundamental to various aspects of software analysis, such as architecture recovery, code smell detection, and software quality evaluation (Mo et al., 2019; Cui et al., 2019; Xiao et al., 2022). Extracting and understanding the dependency interactions between Java and Kotlin code is essential for conducting software analysis in such cross-language software systems. Despite the increasing popularity of Kotlin and its interoperability with Java, there is limited research on how Java and Kotlin interact with each other in real-world projects.
Furthermore, recent studies (Li et al., 2023b) have shown that a large percentage of bug fixes in deep learning frameworks implemented in Python and C++ involve modifying source code in both programming languages. Moreover, the complexity of code changes for such bug fixes is significantly higher than that of fixes in single-programming-language scenarios (Li et al., 2023a). Unlike the interactions between Python and C++, which rely on their compiled public APIs, Java and Kotlin, both being based on the Java Virtual Machine (JVM), can extensively depend on each other at the source code level. This raises the question of whether Kotlin-Java software systems also incur higher maintenance costs and are prone to cross-language bugs, similar to other multi-language systems like Python-C++. Moreover, objects passing through the Python-C interface can incur vulnerabilities due to the different rules in these two languages. For example, if handled incorrectly, a Python list passed to a C program can incur Buffer Overflow due to the fixed-size array in C (Li et al., 2022). We are curious about the challenges developers face in the implementation and maintenance of Kotlin-Java systems and whether similar bugs or vulnerabilities exist in the Kotlin-Java interface. Understanding these challenges can assist researchers and practitioners in designing solutions and tools to navigate these issues and ensure the development of high-quality Kotlin-Java systems.
Although Google Android have provided documentation to guide developers on integrating Kotlin and Java (and, [n. d.]), the challenges faced in real-world projects remain largely unknown. In this paper, we propose to investigate into the dependency relations between Kotlin and Java in real-world projects to better understand the challenges associated with their interactions. To this end, we conducted an empirical study of 23 open-source projects, comprising 3,227 Java and 8,630 Kotlin source files. We developed a static code analysis tool to extract 11 kinds of dependency relations between Kotlin and Java code entities within these projects. Our findings indicate that Java and Kotlin frequently interact in these cross-language projects, with access and call dependency types being the most prevalent.
Compared to files interacting solely within the same language, Java/Kotlin source files that act as the interface for cross-language interactions have undergone more commits, but not necessarily contain a higher volume of lines of code (LOC). Furthermore, we identified seven common mistakes in Kotlin-Java problematic interactions, which deserve special attention from developers. To the best of our knowledge, this work is the first to explore Kotlin-Java dependency relations and to expose common problems in these interactions within real-world projects. The dependency analysis tool that we have implemented and the recurring patterns in Kotlin-Java problematic interactions can assist developers in comprehending and uncovering potential pitfalls in such cross-language systems.
To summarize, our contributions in this paper are as follows:
  • We implemented a static code analysis tool named DependExtractor to extract dependencies between Java and Kotlin, addressing a gap that existing static code analysis tools cannot resolve regarding such dependencies.
  • We demonstrated that the source files involved in Kotlin-Java interactions require undergoing more commits than those only involved in a single language. It also proves that the Kotlin-Java interface is actively changing.
  • We identified seven common mistakes which cause bugs and maintenance effort in Kotlin-Java problematic interactions. These common mistakes can serve as a taxonomy and deserve developers’ special attention.
Data Availability Statement: The data and implemented tool can be accessed through this link (Anonymous, 2024).
2. Approach
2.1. Kotlin-Java Dependency Extraction
DependExtractor Tool: We implemented a tool called DependExtractor in pure Kotlin to extract dependencies between Kotlin and Java code entities. DependExtractor first utilized the Kotlin compiler to resolve Kotlin and Java source files into PSI (Program Structure Interface, similar to Abstract Syntax Tree) elements, including their types, reference types, and locations. Subsequently, it extracted dependency relations between pairs of PSI elements. Finally, these dependencies into a graph containing all PSI elements are aggregated and can be exported in JSON format, with each node representing a Kotlin or Java source file and each edge representing the dependencies between two source files. Following the definitions of dependency relations from related works (Jin et al., 2019; dep, [n. d.]; und, [n. d.]), DependExtractor is capable of extracting 9 types of Java-Java dependencies, 14 types of Kotlin-Kotlin dependencies, and 11 types of Kotlin-Java dependencies. As this paper focuses mainly on Kotlin-Java interactions, we list the 11 types of Kotlin-Java (K-J) or Java-Kotlin (J-K) dependencies and their descriptions in Table 1.
As observed from Table 1, Access signifies that an expression in one source file accesses a field or property of a class in another source file, while Call indicates that an expression in one source file calls a method of a class in another source file. Note that we not only record the dependency type between two source files, but also resolve dependency relations at a fine-grained granularity, capturing details such as the expression statement, locations, and type information for each specific dependency. This fine-grained approach can be valuable for future refactoring research, including but not limited to null safety addressing, exception handling, etc.
Table 1. Supported K-J or J-K Dependency Types
DepType Description
Access an expression accesses a Kotlin/Java field or Kotlin property
Call an expression calls a method
Create an expression creates an object
Extend a class extends a parent class
Extension a Kotlin method or property extension
Local Variable Typed (LVT) a local variable uses a class as its type
Property Typed (PT) a Kotlin property uses a Java class as its type
Implement a class implements an interface
Import a statement imports a type
Parameter a method uses a class as its parameter
Return a method returns a class type
PT and Extension occur from Kotlin entities to Java entities; Other DepTypes can occur either from Kotlin to Java or from Java to Kotlin.
DependExtractor Accuracy Validation: Besides a large amount of unit testing, we also used a Kotlin-Java project with verified dependencies as the ground truth to conduct an integrated test. This ground truth project (implemented in 108 Java and 22 Kotlin source files) is our pre-release version of DependExtractor. Just like software systems which migrate from Java to Kotlin, our tool also went through this process. As we are already familiar with the Kotlin and Java source files in this project, we feel confident to inspect the dependencies of it. In addition, this project contains 5,424 dependency instances, including all 11 dependency types in K-J or J-K, which is a good candidate to test DependExtractor’s ability to extract various dependency instances. Specifically, two authors, each with 4 years of Java and 3 years of Kotlin development experience, traced entities line by line in each source file to list the dependency instances independently. After completing this, they discussed the results until an agreement was reached. This manual inspection served as ground truth and were compared with the JSON result generated by Depends-Kotlin. The comparison shows that our tool can correctly capture 5,387 out of 5,424 dependencies (98.02% precision and 99.34% recall), increasing our confidence in applying DependExtractor in future experiments. Due to space limitations, the inspected ground truth and the results from the DependExtractor tool are shared in the provided Zenodo link (Anonymous, 2024).
2.2. Common Mistake Identification in Kotlin-Java Interaction
To identify Kotlin-Java problematic interactions, we initially mined code repositories to narrow the problematic commit scope for manual inspection. Subsequently, we conducted a thorough manual inspection of code changes to identify and classify code patterns related to Kotlin-Java problematic interactions.
Table 2. Demographic Information of Kotlin-Java Projects in Our Study
ProjectName #Stars Kotlin(%) Java(%) Domain Repository Link
bilimiao2 1300 73.5 24.7 Android https://github.com/10miaomiao/bilimiao2
DataStats 32 81.8 17.9 Android https://github.com/takke/DataStats
dgs-framework 3000 82.2 16 Library https://github.com/Netflix/dgs-framework
FDPClient 353 65.4 34 Minecraft https://github.com/SkidderMC/FDPClient
Flap 286 52.4 47.6 Android https://github.com/AlanCheen/Flap
grpc-kotlin 1100 84.3 7.1 Library https://github.com/grpc/grpc-kotlin
husi 192 74.2 14.2 Android https://github.com/xchacha20-poly1305/husi
intellij-powershell 67 73.1 6.4 IDEA Plugin https://github.com/ant-druha/intellij-powershell
java-stellar-anchor-sdk 34 62.2 37.7 Server backend https://github.com/stellar/java-stellar-anchor-sdk
javalin 7200 89 10.6 Library https://github.com/javalin/javalin
jetbrains 36 74.8 24.4 IDEA Plugin https://github.com/sourcegraph/jetbrains
karibu-testing 104 78.6 6 Library https://github.com/mvysny/karibu-testing
legado 23848 77.6 14.3 Android https://github.com/gedoor/legado
LiquidBounce 1227 81.5 9.9 Minecraft https://github.com/CCBlueX/LiquidBounce
MyExpenses 712 64.8 34.5 Android https://github.com/mtotschnig/MyExpenses
NightX-Client 103 61.3 38.4 Minecraft https://github.com/Aspw-w/NightX-Client
RootEncoder 738 55.8 42.1 Android https://github.com/pedroSG94/RootEncoder
SeriesGuide 1910 68.4 31.6 Android https://github.com/UweTrottmann/SeriesGuide
simbrain 87 60.7 38 Desktop App https://github.com/simbrain/simbrain
SkyHanni 172 84.2 15.7 Minecraft https://github.com/hannibal002/SkyHanni
SkytilsMod 956 90.1 9.7 Minecraft https://github.com/Skytils/SkytilsMod
tasks 3275 92.4 7.5 Android https://github.com/tasks/tasks
wire 4172 62.1 29.4 Library https://github.com/square/wire
2.2.1. Locate Commits to Address Kotlin-Java Problems
We conducted the following filtering process to narrow the scope of potential commits to address Kotlin-Java issues.
  • f1)
    A commit’s comment must contain the words ‘fix,’ ‘bug,’ ‘issue,’ or ‘exception’, and the modified files in a commit must contain file paths with extensions either .java or .kt. All commits that satisfy condition f1 will be added to a collection C1 and proceed to the next filter.
  • f2)
    The modified Java or Kotlin source files in a commit must be involved in Kotlin-Java interactions. For each project, we can obtain the dependency graph of its latest release using the approach outlined in Section 2.1 and identify the files in Kotlin-Java dependency edges. If a commit contains modified files on Kotlin-Java dependency edges, we add it to collection C2 and proceed to the process of manual inspection.
2.2.2. Identify Common Mistakes of Kotlin-Java Problems through Manual Inspection
We establish a protocol for this manual inspection process. First, we located the code changes in Kotlin or Java source file. Then we checked whether the code changes are related with Kotlin-Java interactions. We did this by tracing the code changes down to its reference type. For instance, if a line of code in a Kotlin file is deleted, we get the variables and methods in this line and track the variables down to check whether it is a type of Java class, method, etc. Subsequently, we check whether any new code snippets are added in the Kotlin file near the location of the deleted code line. If present, we compared the differences between the added line and the deleted line. Last, we check the commit comments and discussed what problems developers are trying to address and why developers made such changes.
This step required the dedicated efforts of 2 students with 4 years of Java and 3 years of Kotlin development experience, and one senior researcher with 14 years of Java and 3 years of Kotlin development experience. We followed the practice described in (Campbell et al., 2013), allowing the senior researcher to establish the procedure for identifying code changes and training the students with examples. Then, the two students collaborated to identify code changes and categorize the changes’ reasons. If disagreements occurred, the senior researcher would review the specific cases and discuss them with the two students until a consensus was reached.
3. Study Design
3.1. Subjects
The objective of our study is to investigate the interactions between Kotlin and Java code within real-world projects, and we need to mine open source projects with both Kotlin and Java code. To achieve this, we first applied GitHub API to search repositories with Kotlin code through the GitHub link [^1] https://github.com/search?q=language%3AKotlin+&type=repositories. We further filtered those Kotlin repositories according to the following criteria:
  • 1)
    The repository must contain Java code alongside Kotlin, with Java code constituting more than 3% of the total code;
  • 2)
    The repository should have a star count of no less than 20;
  • 3)
    The repository must have been updated within the last 48 hours relative to our data collection time, which is 12:00 PM on Jan 15, 2024.
Since GitHub does not support searching repositories by multiple languages, we clicked and manually inspected each repository to assess the ratio of programming languages and to fulfill Criterion 1). We established a minimum threshold of 3% for Java code to ensure its active participation in the projects’ functionalities, not just some random Java code. Therefore, we chose a 3% threshold for Java code, rather than 1% or 2%. Criteria 2) and 3) focus on selecting repositories that are actively maintained, which are essential for identifying relevant challenges. By applying these criteria, we identified 23 projects, and their demographic information is presented in Table 2.
These 23 projects span various domains, with 9 focused on Android app development, 5 on Minecraft games, 2 on IntelliJ IDEA Plugins, and 7 encompassing other areas. The percentage of Java code ranges from 6% in the karibu-testing project to 47.6% in the Flap project. In contrast, the percentage of Kotlin code varies from 52.4% in the Flap project to 92.4% in the tasks project. The wide and diverse range of Java and Kotlin code present in these projects enables us to gain a deeper understanding of Kotlin-Java interactions.
3.2. Research Questions
We investigated the following Research Questions (RQs) to gain insight into Kotlin-Java interactions and challenges associated with their interactions.
RQ1. Are dependencies between Kotlin and Java prevalent in Kotlin-Java projects? Kotlin is designed to be fully interoperable with Java, allowing code entities to be seamlessly accessed across both languages. Ideally, dependencies between Kotlin and Java should exhibit the same characteristics as those within homogeneous Kotlin or Java codebases. The answer to this question can enhance our understanding of such cross-language software systems.
RQ2. Do files involved in Kotlin-Java interactions cost more commits and lines of code (LOCs) than files without Kotlin-Java interactions? Recent studies on multi-language systems, such as those involving Python and C++, have shown that, to address bugs occurring in the interface of two languages, developers need to modify code in the interface, resulting in a higher number of lines of code (LOCs) (Li et al., 2023a). This observation raises the question of whether Kotlin-Java software systems are also prone to such cross-language maintenance activities. The answer to this question could pinpoint maintenance hotspots within these Kotlin-Java systems.
RQ3. Are there common mistakes in Kotlin-Java interactions? Previous research proposed various techniques for detecting code smells or anti-patterns within a single-language codebase. More recent studies have found that bugs and security problems in multiple language systems (Li et al., 2023b). For instance, a recent study has uncovered security issues arising from interactions between Java and C code (Li et al., 2022). By investigating whether there exist common mistakes in Kotlin-Java interactions, we aim to evaluate the problems associated with Kotlin-Java interactions and pinpoint the scenarios that may require closer attention when handling such interactions.
4. Results
4.1. Dependencies in Kotlin-Java Projects
We applied DependExtractor on the most recent versions of our 23 selected projects, and the results are presented in Table 3. It is interesting that in these 23 projects we did not observe any Java-Kotlin dependencies (DependExtractor was tested thoroughly). This may be due to our subject selection criteria, and the primary language of these projects is Kotlin, in which developers tend to use Kotlin to depend on Java for functionalities. All dependencies in this table are counted without weights. The last column shows the ratio of Kotlin-Java dependencies to the minimum of Java-Java and Kotlin-Kotlin dependencies. Similar to the Barrel principle, if a project contains only a few Java source files, the number of cross-language dependencies should not be disproportionately large. Therefore, the smaller number of source files determines the cross-language dependencies. We compared Kotlin-Java dependencies with the minimum of J-J and K-K to check whether cross-language dependencies are more prevalent than single-language dependencies.
Table 3. Overall Dependencies in the Subjects
ProjName #J #K J-J K-K K-J $\frac{K-J}{Min(J-J,K-K)}$%
bilimiao2 88 489 222 5979 122 55.0
DataStats 9 10 2 24 46 2300.0
dgs-framework 108 305 93 1648 567 609.7
FDPClient 242 621 320 8886 1291 403.4
Flap 20 104 19 636 19 100.0
grpc-kotlin 8 83 2 285 8 400.0
husi 29 175 81 1913 494 609.9
intellij-..shell 7 162 6 929 11 183.3
java..-sdk 548 260 1381 810 4796 592.1
javalin 57 260 46 2600 635 1380.4
jetbrains 95 300 94 2104 204 217.0
karibu-testing 7 112 6 171 3 50.0
legado 87 746 269 12057 89 33.1
LiquidBounce 84 575 11 9114 11 100.0
MyExpenses 325 703 678 9756 2842 419.2
NightX-Client 184 253 286 3223 1244 435.0
RootEncoder 136 334 340 3209 395 116.2
SeriesGuide 176 391 216 4483 1468 679.6
simbrain 325 481 756 8140 1434 189.7
SkyHanni 302 697 301 14271 1898 630.6
SkytilsMod 89 302 8 3639 107 1337.5
tasks 83 639 63 9028 1245 1976.2
wire 218 628 168 9935 60 35.7
Average 528.6
#J and #K denote the number of Java and Kotlin files; J-J, K-K and K-J denote the dependencies number from Java to Java, Kotlin to Kotlin, and Kotlin to Java entities, respectively.
As shown in Table 3, Kotlin and Java source files frequently interact. For instance, in the bilimiao2 project, there are 122 Kotlin-Java dependencies within 88 Java and 489 Kotlin files. If we divide 122 Kotlin-Java dependencies by the minimum of Java-Java and Kotlin-Kotlin dependencies, we obtain 55.0%. This significant ratio indicates that Kotlin-Java dependencies are prevalent in this project. Similar results are observed in other projects. In the DataStats project, for example, there are 46 Kotlin-Java dependencies within 9 Java and 10 Kotlin files, outnumbering 2 Java-Java and 24 Kotlin-Kotlin dependencies. The resulting ratio of 2300.0% underscores the dominance of Kotlin-Java dependencies in this project. Overall, the last column of this table reveals that the ratio of Kotlin-Java dependencies ranges from 33.1% in the legado project to 2300.0% in the DataStats project, with an average of 528.6%. This average indicates that, on average, Kotlin-Java interactions are five times more prevalent than the minority of single-language interactions.
Figure 1. Circular Layout of Dependencies
Table 4. Statistics of Dependencies from Kotlin to Java entities (K-J) in the subjects
ProjName Access Call Create Extend Extension LVT PT Implement Import Parameter Return
bilimiao2 8 43 3 1 1 25 41 3 39 15 9
DataStats 163 89 0 0 0 29 13 0 9 0 0
dgs-framework 153 112 953 8 0 66 14 4 187 50 6
FDPClient 5919 729 31 54 0 846 99 0 276 12 16
Flap 14 3 3 0 0 7 4 0 3 0 1
grpc-kotlin 89 3 1 0 0 0 6 0 0 1 0
husi 1523 81 120 14 64 167 4 0 111 40 64
intellij-powershell 8 3 1 0 0 1 0 0 3 0 1
java-stellar-anchor-sdk 5382 5115 2126 20 1 3462 882 3 1843 100 2752
javalin 204 1315 29 0 2 157 69 4 307 90 745
jetbrains 90 68 26 4 0 55 35 1 69 15 25
karibu-testing 1 0 1 0 0 0 0 0 0 1 0
legado 5 51 32 0 0 21 14 4 24 18 8
LiquidBounce 8 9 0 0 0 0 13 0 3 0 0
MyExpenses 3623 1365 124 63 5 1420 1086 40 2646 243 737
NightX-Client 4002 1277 141 40 0 443 141 0 333 3 23
RootEncoder 243 399 65 6 0 46 33 9 134 47 6
SeriesGuide 1142 581 700 94 4 488 346 23 506 147 43
simbrain 522 711 321 36 14 402 286 14 411 110 124
SkyHanni 2042 24 16 98 21 519 887 1 280 71 294
SkytilsMod 83 20 0 0 0 19 7 0 49 1 1
tasks 378 874 520 4 1 398 354 8 460 87 77
wire 45 76 18 4 0 50 7 0 24 0 4
Total 25647 12948 5231 446 113 8621 4341 114 7717 1051 4936
Figure 1 depicts the circular layout of dependencies in three projects. The nodes on the circle’s edge represent source files. The blue, red, and purple curved lines connecting two nodes represent Kotlin-Kotlin, Java-Java, and Kotlin-Java dependencies, respectively. It is evident that in these three projects, Kotlin-Java dependencies are not confined to a specific area; instead, they are spread across the entire systems. We observed this in 17 out of 23 projects, except for the Flap, grpc-kotlin, intellij-kotlin, karibu-testing, legado, and LiquidBounce projects. One explanation is that the migration process, legacy code, or design in these systems differs from the other 17 systems. We will investigate these observations in the future.
Table 4 presents the statistics of Kotlin-Java dependency types in each project. As observed in this table, certain dependency types are present in some projects but not in others. For instance, the “Extension” dependency type is found in 9 projects, while the remaining 14 projects do not exhibit this type. We suspect this discrepancy is due to the distinct characteristics of the projects. However, all 11 defined types of Kotlin-Java dependencies can be observed across all projects. The top two most frequent dependency types among Kotlin-Java interactions are “Access” and “Call”.
To answer RQ1, the dependency extraction from 23 open-source Kotlin-Java projects spanning various domains reveals that Kotlin-Java interactions are widespread, thus validating the interoperability between Kotlin and Java. Kotlin-Java dependencies can occur more frequently than minor single-language interactions, with “Access” and “Call” being the top two most frequent dependency types. Furthermore, in most projects Kotlin-Java dependencies are distributed across the entire system rather than confined to a local area.
Table 5. Maintenance Cost of Kotlin/Java Source Files in Kotlin-Kotlin, Java-Java, and Kotlin-Java
ProjectName Kotlin Files in Kotlin-Kotlin Java Files in Java-Java Kotlin Files in Kotlin-Java
#Avg_Cmt_K Avg_LOC_K #Avg_Cmt_J Avg_LOC_J #Avg_Cmt_Kj Avg_LOC_Kj
bilimiao2 17 116.5 1.5 5.7 2.8 49.7
DataStats 4.6 22.0 4.5 47.0 20.5(4.5x) 27.2
dgs-framework 10.8 97.0 3.3 22.0 6.3 81.1
FDPClient 11.1 60.6 14.8 110.1 37.5(3.4x) 53.9
Flap 13.3 102.1 12 58.6 27.9(2.1x) 61.8
grpc-kotlin 6.5 30.0 5 14 11(1.7x) 32.3
husi 3.9 24.0 3.5 34.8 23.7(6.1x) 20.8
intellij-powershell 11.5 53.7 11.3 90.3 38.3(3.3x) 26.3
java-stellar-anchor-sdk 7.3 113.9 7.3 52.9 14.9(2.0x) 102.3
javalin 13.9 78.3 5.6 23.4 8.6 62.5
jetbrains 13.2 78.2 10.3 76.9 36.7(2.8x) 44.6
karibu-testing 10.5 99.5 8.4 62.6 3.4 121.7
legado 41.9 825.3 2.8 8.5 3.7 170.3
LiquidBounce 9 62.5 24 182 49(5.4x) 47.8
MyExpenses 18 102.6 42.7 228.1 103.3(5.7x) 84.7
NightX-Client 6.4 34.2 7 46.3 23.1(3.6x) 33.2
RootEncoder 15.6 90.5 14.9 87.0 26.8(1.7x) 46.0
SeriesGuide 7.3 41.6 7.4 37.7 17.5(2.4x) 40.0
simbrain 10.9 86.3 8.5 59.6 31.4(2.9x) 57.7
SkyHanni 21.9 117.1 6.5 42.1 17.2 90.3
SkytilsMod 17.9 102.9 10.3 194.5 31.2(1.7x) 35.6
tasks 11.7 65.4 11.1 66.1 35.6(3.0x) 49.1
wire 1.6 8.5 1.7 9.3 3.8(2.4x) 39.4
#Avg_Cmt_K and Avg_LOC_K denote the average commit number and LOC for Kotlin files only involving K-K dependencies;
#Avg_Cmt_J and Avg_LOC_J denote the average commit number and LOC for Kotlin files only involving J-J dependencies;
#Avg_Cmt_Kj and Avg_LOC_Kj denote the average commit number and LOC for Kotlin files participating K-J dependencies.
4.2. Maintenance Cost in Kotlin-Java Projects
Recent studies on multi-language systems have revealed that the interface between two languages is vulnerable and susceptible to bugs, often resulting in increased maintenance costs for fixing these issues (Li et al., 2023a, 2022, b). This raises the question of whether the interface between Kotlin and Java faces similar challenges.
This research question aims to gain a better understanding of Kotlin-Java interactions. For each project, utilizing the dependency information got in RQ1 (see Section 4.1), we extracted the paths of Kotlin or Java source files involved in Kotlin-Java dependencies, denoted as Paths(kt/java). Subsequently, we traversed each project’s git commits and matched each modified file path of a commit with Paths(kt/java). If a match was found, we incremented #commit(aFile) by 1 and added the modified lines of code (including insertions and deletions) to #LOC(aFile). Similarly, we calculated #commit and #LOC for Java or Kotlin source files participating only in Java-Java or Kotlin-Kotlin interactions.
Table 5 presents the average LOC and commits for different types of files in each project. The first two columns denote the average commit counts and LOC for Kotlin source files that only interact with other Kotlin source files. The middle two columns denote the average commit counts and LOC for Java source files that only interact with other Java source files. The last two columns denote the average commit counts and LOC for Kotlin source files that interact with Java source files.
Comparing the first column with the second-to-last column, we observe that in 17 out of 23 projects (highlighted in orange), the average number of commits for Kotlin source files participating in Kotlin-Java interactions is significantly larger than for Kotlin source files not participating in Kotlin-Java interactions. The number in the parentheses denotes the ratio of these two commit counts. For example, in the husi project, the average number of commits (23.7) for Kotlin source files participating in Kotlin-Java interactions is 6.1 times of that (3.9) for Kotlin source files not participating in Kotlin-Java interactions. On average across all 23 projects, Kotlin source files participating in Kotlin-Java interactions have 2.5 times more commits than Kotlin source files not participating in Kotlin-Java interactions.
However, as observed from Table 5, the LOC for Kotlin source files participating in Kotlin-Java interactions show no significant difference from Kotlin source files not participating in Kotlin-Java interactions. Only in these four projects (highlighted in orange in the last column) Kotlin source files participating in Kotlin-Java interactions have more LOCs on average. Furthermore, we conducted a t-test for these columns, and the t-test results between column #Avg_Cmt_K and #Avg_Cmt_Kj indicate a statistically significant difference at a 95% confidence level (df = 22, p-value = 0.016). Such results support the alternative hypothesis that #Avg_Cmt_Kj is significantly greater than #Avg_Cmt_K. However, the t-test results indicate no statistically significant difference between #Avg_LOC_K and #Avg_LOC_Kj (df = 22, p-value = 0.125). We conjecture this is because Kotlin source files participating in the interactions are often committed to modify only a small amount of code related to the facade, such as method calls, field access, parameter changes.
To answer RQ2, the average number of commits and LOC for 23 open-source Kotlin-Java projects indicate that Kotlin source files participating in Kotlin-Java interactions tend to have more commits but not more LOCs.
4.3. Common Mistakes in Kotlin-Java Interactions
The results of RQ2 reveal that source files participating in Kotlin-Java interactions undergo more commits and are prone to changes. To answer RQ3, we delve into the details of why certain Kotlin-Java interactions are problematic. Following the approach described in Section 2.2, we identified common mistakes in Kotlin-Java problematic interactions.
Table 6. Locate Problematic Kotlin-Java Commits
All commits 83792
f1: A commit’s comment contains words ‘fix’, ‘bug’, ‘issue’ or ‘exception’ 5137
f2: A commit’s modified files are participated in Kotlin-Java interactions 1314
Manual Inspection: confirmed commits to fix problematic Kotlin-Java interactions 103
Table 6 presents the data in the process of locating commits involving problematic Kotlin-Java interactions. Initially, there were 83,792 commits among these projects. After filtering commits with comments containing the words ‘fix’, ‘bug’, ‘issue’ or ‘exception’, there were 5,137 remaining commits. These 5,137 commits were compared with the dependency graphs of the latest version, resulting in 1,314 commits with modified files participating in Kotlin-Java interactions. Finally, we manually inspected these 1,314 commits and confirmed 103 commits containing fixes for problematic Kotlin-Java interactions. After discussion among three developers, we classified 98 of these 103 problematic Kotlin-Java interactions into 7 categories according to the issue and fixing strategies. Table 7 lists these categories and their occurrences. In the following, we provide an example for each of these 7 categories.
Table 7. Common Mistakes in Kotlin-Java Interactions
Common Mistake Category Occurrence
Java Return Value NPE 47
Java Invoke NPE 21
Kotlin Wrong Not Null Assertion 12
Call Chain NPE 9
Kotlin Inherit Java Nullability Inconsistency 5
Immutable Collection Conversion 3
Overuse Java Optional in Kotlin code 1
Other 5
Total 103
4.3.1. Java Return Value NPE (NullPointerException): fix by adding Null-safe Operator for returned values from Java
This is the most common mistake we observed among the 103 problematic interactions. When Kotlin code calls a Java method, the Kotlin compiler will not be able to determine the nullability of the return value if the Java method does not indicate the nullability of the return value with annotations such as @NotNull or @Nullable. If Kotlin code accesses a variable whose nullability is unknown, a NullPointerException (NPE) may get triggered.
Figure 2. Kotlin Add Null-Safe Operator to Fix NPE
Figure 2 shows an example of this mistake in commit 2cd4ca1222https://github.com/hannibal002/SkyHanni/commit/2cd4ca1 in the SkyHanni project. The code marked in red attempts to call the method getCanonicalName() in the Java class ‘Class’ and assumes that the returned value from this method is not NULL. However, at runtime this method returns a NULL and the red line throws a NPE. As stated in the commit comment, the commit aims to “fix an NPE in ReflectionUtils.shPackageName”. It added the Null-Safe Operator ? to the returned value of the getCanonicalName() method. In this way, the Null-Safe Operator ? performs a NULL check on the returned value shown in the lime color. If it is NULL, it stops the process without proceeding to the next call.
4.3.2. Java Invoke NPE: fix by clarifying Nullability on Java code
This issue is similar to the previous one but differs in the error location and fixing strategy. When developing a Kotlin-Java system, adding appropriate nullable or non-nullable annotations (@NotNull or @Nullable) to Java’s public API will enable modern IDEs to alert developers its nullability at the time of invocation and provide stricter code checking during compilation. Without annotations on parameters of Java’s public API, Kotlin code is unable to determine the required nullability of parameters and an inappropriate API call will result in a NPE.
Figure 3 shows commit f240ed6333https://github.com/ant-druha/intellij-powershell/commit/f240ed6 from the intellij-powershell project, which is a good example of the importance of Java nullability annotation when interacting with Kotlin. To “fix exception thrown when plugin is enabled but no powershell is found in the path” as stated in the commit comment, @NotNull annotation was added to the powerShellExecPath Java method (shown in lime color) and Kotlin code which invoked this method was changed accordingly.
Figure 3. Java API Add Annotation to Fix NPE
4.3.3. Kotlin Wrong Not Null Assertion: fix by modifying non null assertion operator
When a developer is confident that the returned value of a method or a property is non-null at all times, they apply a not null assertion !! operator in Kotlin code to emphasize its non-null nature. Such assertions offer significant convenience in subsequent code writing, eliminating the need for NULL checks and enhancing logical flow. However, not-null assertions also pose risks. If applied to a value whose nullability is not entirely determinable, an NPE may still occur at runtime if the value turns out to be null, potentially affecting a substantial portion of the subsequent code.
Figure 4. Kotlin Not Null Assertion throws NPE
Figure 4 depicts commit 83dc35e [^4] https://github.com/square/wire/commit/83dc35e from the wire project. A not null assertion !! operator was placed on the entries variable. However, this code throws a NPE at runtime because it fails to fully consider the nullability of the entries variable. The commit message “fix: options entries can be null for files imported from external jars” verified this observation and the not null assertion !! operator was changed to the null-safe Operator ? in this commit.
4.3.4. Call Chain NPE: fix by adding null safe operation or add annotation on method signature
In interactions between Kotlin and Java, we frequently observe that Kotlin code performs method chain calls on Java methods. If a method in the sequence returns a null value and the subsequent method call relies on this result without checking, it can lead to an NPE. The commit 79c9792555https://github.com/ant-druha/intellij-powershell/commit/79c9792 in the intellij-powershell project demonstrates an example of this case.
1/* SyntheticPowerShellCodeBlock.kt */
2import com.intellij.lang.ASTNode
3import com.intellij.psi.formatter.java.LeafBlock
4
5 private fun isRParenth(block: Block?): Boolean {
6 return block is LeafBlock && block.node.elementType $=$ PowerShellTypes.RP
7 }
In this example, a method in the SyntheticPowerShellCodeBlock Kotlin Class is importing and depending on Java classes ASTNode and LeafBlock. The code in Line 6 shows a calling chain to LeafBlock and ASTNode through block.node.elementType in this method. If the node attribute is not initiated in block object, the subsequent call .elementType will throw a NPE. In this commit, developers patched this bug by modifying it to block.node?.elementType. The null safe operator ? in block.node? performs a NULL check on the node object and allow for a safe call. If it is a null, block.node?.elementType will return a null without calling .elementType.
There are two ways to fix this issue. One is to use null safe operator just like in this example to allow the returned value being nullable. The other is to specify the return value of each method nullable by adding annotation along with the method’s signature and let IDEs handle it at the compilcation.
4.3.5. Kotlin Inherit Java Nullability Inconsistency: fix by ensuring parent-child API consistent
This scenario happens when a Kotlin class inherits a Java class and overrides the parent methods, the nullability of parameters in the overridden methods is inconsistent between the parent and child class. This is not a problem in Kotlin-Kotlin or Java-Java interactions as the type consistency is manually enforced in modern IDEs. However, the nullability consistency is not enforced for Kotlin-Java interactions by modern IDEs so developers need to manually check it. The commit da56686666https://github.com/grpc/grpc-kotlin/commit/da56686 in the grpc-kotlin project shows an example and how developers were addressing it.
As we can see from the following code, DeclarationSubject class is written in Kotlin and it extends the Subject Java class. The second parameter Declarations in the child Kotlin class is not null while the parent Java class Subject specify the second parameter actual as @Nullable. There exists an inconsistency between these two constructor methods, though this does not cause any problems before version 1.1.3 as the callers of these constructors in the Factory interface does not specify @Nullable in its parameter.
1 /** Subject.java */
2 protected Subject(FailureMetadata metadata, @Nullable Object actual) {
3 this(metadata, actual, /* typeDescriptionOverride= */ null);
4 }
1 /** DeclarationsSubject.kt */
2 class DeclarationsSubject(failureMetadata: FailureMetadata, private val actual: Declarations): Subject(failureMetadata, actual) {...}
3 // This calls the Factory interface in Subject.
4 val declarationsSubjectFactory: Subject.Factory<DeclarationsSubject, Declarations> =
5 Subject.Factory(::DeclarationsSubject)
However, in Version 1.1.5, the Factory interface in Subject class added @Nullable annotations to its parameter ActualT actual as shown below. When the grpc-kotlin project upgraded the Subject’s package – com.google.truth:truth–, from Version 1.1.3 to 1.1.5, the Factory interface requires a nullable type for createSubject’s second parameter specifically while DeclarationSubject is not. This results in a complication error.
1 /** Factory in Subject.java in version 1.1.3. */
2 public interface Factory<SubjectT extends Subject, ActualT> {
3 /** Creates a new {@link Subject}. */
4 SubjectT createSubject(FailureMetadata metadata, ActualT actual);
5 }
6 /** Factory in Subject.java in version 1.1.5. */
7 public interface Factory<SubjectT extends Subject, ActualT> {
8 /** Creates a new {@link Subject}. */
9 SubjectT createSubject(FailureMetadata metadata, @Nullable ActualT actual);
10 }
To fix this error, developers added the null safe operator ? in the child Kotlin class DeclarationSubject, allowing the second parameter to be nullable. From this example, we can conclude that, although the inconsistency would not result in compilation errors within the DeclarationSubject class itself, it can arise potential issues from a third class’ calling. It is necessary to make sure that Kotlin-Java parent-child overridden methods strictly follow the same nullability type. Furthermore, this is also a violation of Liskov Substitution Principle, which requires a subtype should be substitutable for its supertype.
4.3.6. Immutable Collection Conversion: fix by creating a defensive copy of Java collections
In Kotlin, collections can be categorized as modifiable, such as MutableList or MutableSet, and unmodifiable types, such as List or Set. If Java code calls collections from Kotlin code, modern IDEs ensure that proper methods of collections are invoked, meaning that mutation methods like add() or remove() in Java cannot be applied to an unmodifiable Kotlin collection.
However, if Kotlin code is accessing collections from Java code, the mutability of the collections is unknown to the Kotlin side. This is because unmodifiable collections in Java act as a wrapper around existing modifiable collections. Therefore, public methods such as add, remove, and update in Java modifiable collections can also be called on Java unmodifiable collections. This presents a challenge on the Kotlin side because the mutability of collections returned from Java is not known.
Figure 5 illustrates an example of this issue in the commit a8cf6eb [^7] https://github.com/square/wire/commit/a8cf6eb of the wire project. The getFromBuilder method utilizes Java’s reflection mechanism and returns an object with a java.util.Map type. Prior to this commit, this Map object was mistakenly treated as a Kotlin MutableMap. Since the mutability of Java’s collections is not inherently recognized in Kotlin, developers introduced a check for this Map object in this commit. If the object is confirmed to be a MutableMap, records are added to it directly. If it is not a MutableMap, a defensive copy of the map is created first, and then the records are put into the copied map. If the object cannot be cast to either a MutableMap or a Map, an exception is thrown. Although this approach solves the problem, we do not recommend using obj is MutableMap in JVM to check its mutability. The good practice is to assume that all collections of unknown mutability are immutable and create a defensive copy (you, [n. d.]).
Figure 5. Immutable Collection Conversion
4.3.7. Overuse Java Optional in Kotlin code: fix by leveraging safe check mechanisum in Kotlin
One key feature of Kotlin is to eliminate the danger of null references by leveraging nullable types and non-nullable type. Compared to Kotlin, Java does not have such feature and thus it needs other techniques such as Optional to check whether a value is null or not. Optional is a container class introduced in Java 8 and provides a way to represent optional values without using null references, thereby helping to avoid common programming pitfalls associated with null values. The Optional class includes methods like isPresent(), get(), ifPresent(), and orElse(), which enable developers to check value presence and absence and to handle cases where a value might be missing.
We observed a case that Kotlin code overuse Java Optional, instead of leveraging Kotlin’s own typing system. The commit b9853c6888https://github.com/Netflix/dgs-framework/commit/b9853c6 in the dgs-framework demonstrates an example, with its commit comment “eliminate some unnecessary usage of Optional”.
Figure 6. Overuse Java Optional in Kotlin Code
Figure 6 illustrates a portion of the code modifications made to the DgsGraphQLMetricsInstrumentation Kotlin file. Previously, the Optional class was extensively utilized to encapsulate four attributes: queryComplexity, operation, operationName, and querySignature. Within the tags method, the map() and orElse() methods from the Optional class were used to check the nullability of the wrapped objects. However, in this commit, developers opted to removed all this Optional wrapping and leveraged Kotlin’s null mechanisms. For instance, as shown in the last two lines of Figure 6, the null safe operator ? was appended to querySignature to verify its nullability, thereby replacing the need for the map() method from the Optional class. Overuse of the Optional class not only consumes additional memory and can negatively affect system performance, but it may also complicate code comprehension. We strongly suggest practitioners to examine the overuse of Optional in Kotlin code.
To answer RQ3, we conducted a manual inspection of 103 cases involving Kotlin-Java interactions, which had already led to system errors or incurred maintenance costs. Through these inspections, we identified 7 categories of common mistakes along with corresponding fixing strategies. We believe that these categorized cases can serve as a taxonomy and assist developers in correctly handling Kotlin-Java interactions.
5. Discussions
5.1. Analysis of Results
The findings of RQ1 have demonstrated the interoperability between Kotlin and Java. Unlike other cross-language systems that rely on compiled code, the dependencies between Kotlin and Java code are more direct and span across entire systems. Consequently, this leads to more cross-language interfaces and presents greater challenges for Kotlin-Java interactions. The results of RQ2 indicate that source files involved in Kotlin-Java interactions undergo more commits than those exclusively involved in a single language. It demonstrates that Kotlin-Java interfaces are inherently more active, just like other cross-language interfaces, even though Kotlin and Java can coexist within the same IDEs and their interaction mechanism is different from other cross-language systems.
The manual inspections of 103 cases in RQ3 reveal specific 7 categories of mistakes within Kotlin-Java interactions. In contrast to Python-C interfaces, which are known to suffer from vulnerabilities (Li et al., 2022), 89 out of 103 problematic interactions between Kotlin and Java are related to NullPointerExceptions (NPEs). Despite Kotlin being known for its better null safety mechanism compared to Java, the interface between Kotlin and Java still experiences NPEs. Other mistakes, such as “Immutable Collection Conversion” and “Overuse of Java Optional in Kotlin”, result from unfamiliarity with Kotlin documentation. It is also surprising that we did not find any exception handling mistakes in the Kotlin-Java interactions, despite Java and Kotlin having different exception catching rules. This can be attributed to modern IDEs’s checking in the compilation process.
5.2. Implications
To Software Practitioner: Although Kotlin and Java can interact smoothly at the code level, common interaction mistakes still exist that may not be captured by modern IDEs. The identified categories of interaction mistakes and their fixing strategies can serve as a valuable reference, complementing the Kotlin-Java interoperability guide from Android (and, [n. d.]).
To Researchers: Different cross-language interactions may suffer from various problems. Conducting empirical studies to identify problematic interactions with associated errors and maintenance costs is key to pinpointing them accurately. For Kotlin-Java interactions, there is a high demand for research and tools that can automatically detect and fix these issues.
5.3. Threats to Validity
The first threat is the accuracy of our dependency extraction tool. DependExtractor is based on the Kotlin compiler and dependency types are defined according to previous research. Though we built the ground truth of a Kotlin-Java project and validated our tool can achieve a relatively high precision and recall, we cannot guarantee that it is 100% accurate. We will continue to validate and improve our tool’s accuracy in the future.
Another threat is that we associated #commit and LOC with file paths. A file involved in K-J dependencies may undergo activities unrelated to fixing interaction problems, or these activities may be mixed in a commit. We acknowledge that #commit and LOC may not capture the exact maintenance cost of a file. However, there is no practical way to split the cost and directly measure the exact amount. We hope that the qualitative analysis detailed in Section 4.3 can partly mitigate this threat.
An external validity is that we only investigated 23 real-world projects with Kotlin as the primary languages. It is possible that additional categories of problematic Kotlin-Java or Java-Kotlin interactions may be discovered in other projects. Our future work will aim to include more software systems and industry examples to further validate our findings.
6. Related Work
In this section, we discuss related research to our work from two perspectives: empirical studies on Kotlin-Java projects and code dependency extraction in cross-language systems.
6.1. Empirical Studies on Kotlin-Java Projects
Since the introduction of Kotlin, various empirical studies have been conducted to understand Kotlin-Java projects. Martinez and Mateus investigated why developers migrate Android applications from Java to Kotlin (Martinez and Mateus, 2021). They also examined the adoption, usage, and evolution of Kotlin features in Android development (Mateus and Martinez, 2020). Nakamura et al. conducted a performance study of Kotlin and Java programs using bytecode analysis (Nakamura et al., 2024). Ardito et al. compared the effectiveness of Kotlin versus Java in Android app development tasks (Ardito et al., 2020). Flauzino et al. compared code smells between Java and Kotlin (Flauzino et al., 2018). Bose et al. compared Java versus Kotlin in terms of null safety, exception handling, and other aspects (Bose et al., 2018).
While previous studies have compared various characteristics of Java with Kotlin, our study offers an empirical investigation into how Kotlin actually interacts with Java code. Despite the availability of documentation from Google and Android guiding Kotlin-Java interaction, to our knowledge, our study is the first to examine Kotlin-Java interactions in real-world projects. Our findings reveal that while Kotlin and Java demonstrate great interoperability and interact with each other extensively across systems, challenges and issues persist in such interactions. The recurring patterns identified in problematic Kotlin-Java interactions in this paper deserve special attention from developers, especially Kotlin developers.
6.2. Code Dependencies Extraction
The dependency relations among entities in the source code form the foundations of software analysis (Mo et al., 2019; Cui et al., 2019; Xiao et al., 2022). Existing techniques often rely on static code analysis to capture and model extract dependency relations from source code such as Java, Python and C/C++  (Jin et al., 2024; Filgueira and Garijo, 2022; Molina et al., 2021; Savić et al., 2014). Various tools such as Structure 101 (str, [n. d.]), Understand (und, [n. d.]), and DV8 (dv8, [n. d.]) have been developed to support the extraction of dependency relations among entities in large-scale software projects.
While techniques and tools for extracting dependencies among code entities within single programming language are relatively mature, the extraction of dependencies among multiple programming languages remains largely unexplored. With advancements in modern platforms, it has become common to utilize multiple programming languages within the same system to achieve enhanced performance and scalability (Mayer et al., 2017; Tomassetti and Torchiano, 2014). However, this also increases the complexity of systems, posing challenges for extracting dependencies in such heterogeneous langauge environments. A recent systematic review reveals that traditional code analysis tools face challenges in these cross-language systems (Latif et al., 2023). A study also highlights high-risk vulnerabilities between Python and C interfaces, proposing symbolic dependence analysis to detect such vulnerabilities (Li et al., 2022). As a complement to the above studies, our study proposed to extract code dependencies between Java and Kotlin code and further detect problems in Kotlin and Java interactions.
7. Conclusions
In this paper, we conducted an empirical study of Kotlin-Java interactions in 23 real-world projects. Our results demonstrate that Kotlin and Java interact frequently, confirming the extensive interoperability of Kotlin and Java. However, our study also reveals that source files involved in such cross-language systems have more commits compared to those involved in only a single language.
Furthermore, we identified seven common mistakes in Kotlin-Java interactions, along with their corresponding fixing strategies. Unlike code smells, which are detected by rules and often have high false positives, these mistakes were extracted from commits in Git repositories and have been shown to require more maintenance effort. We believe that recognizing these identified mistakes in Kotlin-Java interactions can assist developers in effectively improving such cross-language interactions.
Our future work will focus on the following: first, investigating more systems and industry examples to further validate our findings; second, extending this empirical study to an automatic tool that can detect and fix problematic Kotlin-Java interactions; and third, enhancing our DependExtractor tool to support more cross-language dependencies.